What is the pKb of ClO2-? Hints: What is the Ka of the conjugate acid of ClO2-? You're finding Ka then calculating pKb from that.
Answer #1

Conjugate acid of ClO2 - ion is HClO2

I Conjugate So, ka of acid HUO₂ of clo? ion = 1.18 10-2 is HOO2. We know ka X kb - 10-14 1.1X1024 Kb = 10-14 ko = 10-14 1.1 X